OpenStudy (anonymous):

[9.02] What are the solutions of 6x2 − x − 2 = 0?

OpenStudy (anonymous):

fined the like turms

OpenStudy (anonymous):

1t step : \[6x ^{2} + 3x - 4x - 2 = 0\]

OpenStudy (anonymous):

2nd : \[3x(2x+1) - 2(2x+1) = 0\]

OpenStudy (anonymous):

do you get it so far?

OpenStudy (anonymous):

then the 3rd step would be : (3x-2)(2x+1) = 0

OpenStudy (anonymous):

therefore the solution should be x = \[2\over3\] and \[-1\over2\]

OpenStudy (anonymous):

I hope I am right :) :) happy new years :p